# Gradient-Direction-Aware Density Control for 3D Gaussian Splatting

### v1

@@ 3.3.2 Nonlinear dynamic weight design
Plenoxels fridovich2022plenoxels 0.626 23.08 0.463 2.1GB 0.719 21.08 0.379 2.3GB 0.795 23.06 0.510 2.7GB
INGP muller2022instant 0.671 25.30 0.371 13MB 0.723 21.72 0.330 13MB 0.797 23.62 0.423 13MB
Mip-NeRF360 barron2022mip 0.792 27.69 0.237 8.6MB 0.759 22.22 0.257 8.6MB 0.901 29.40 0.245 8.6MB
3DGS kerbl20233d 0.815 27.21 0.214 734MB 0.841 23.14 0.183 411MB 0.903 29.41 0.243 676MB
Pixel-GS* zhang2024pixelgs 0.823 27.67 0.193 1.2GB 0.853 23.74 0.150 1.05GB 0.896 28.91 0.248 1.1GB
AbsGS-0004 ye2024absgs 0.820 27.49 0.191 728MB 0.853 23.73 0.162 304MB 0.902 29.67 0.236  [^4] MB
AbsGS-0008 ye2024absgs 0.815 27.41 0.211 450MB 0.844 23.54 0.183 202MB 0.903 29.69 0.241 380MB
GDAGS (Ours) 0.821 27.55 0.199 515MB 0.847 23.55 0.176 226MB 0.905 29.70 0.235 388MB
4 Experiments
4.1 Setup
Datasets. Similar to the 3DGS dataset, we selected all 9 scenes introduced in MIP-NERF360 barron2022mip , including indoor and outdoor scenes, Train and Trck scenes in Tanks&Temples hedman2018deep , and drJohnson and playroom scenes in Deep Blending knapitsch2017tanks .
Baselines. The proposed method is benchmarked against three categories of approaches: fast NeRF frameworks (Plenoxels fridovich2022plenoxels , Instant-NGP muller2022instant ), state-of-the-art NeRF-based models (Mip-NeRF360 barron2022mip ), and 3DGS variants including the widely adopted 3DGS kerbl20233d , over-reconstruction-focused AbsGS ye2024absgs , and Pixel-GS zhang2024pixelgs , ensuring comprehensive evaluation across all datasets.
Implementation Details. Our experimental setup follows the original paper parameter settings of 3DGS, with densification performed every 100 iterations, stopping densification after 15k iterations and completing training at 30k iterations. The suppression factor $\alpha$ in the weight function is set to 0.8, the amplification factor $\beta$ is set to the size threshold of 25, and the hyperparameter $p$ is set to 15. All experiments are conducted on a single NVIDIA 4090 GPU with 24GB of graphics memory. We report the view metrics PSNR, SSIM, and LPIPS zhang2018unreasonable , and the memory used to store Gaussian parameters to demonstrate the trade-off between performance and efficiency.
@@ 4.2 Comparison
GDAGS-S 0.819 27.52 0.240 441MB 0.846 23.54 0.178 195MB 0.905 29.67 0.238 328MB
GDAGS-C 0.812 27.46 0.217 615MB 0.847 23.71 0.180 305MB 0.904 29.70 0.240 452MB
GDAGS (Ours) 0.821 27.55 0.199 515MB 0.847 23.55 0.176 226MB 0.905 29.70 0.235 388MB
Quantitative Results. We report the quantitative results of the experiment in the Table 1, which indicate that our method produces comparable results in most cases. It is worth noting that our method has less memory at performance close to or better than SOTA. Compared to Pixel-GS, GDAGS achieves comparable or superior rendering performance with only 20%-50% of the memory consumption. This, combined with its comparison against 3DGS, demonstrates that GDAGS effectively addresses over-densification through nonlinear dynamic weights that suppress redundant Gaussians during densification. By adaptively controlling density, GDAGS enables scene Gaussians to converge toward compact spatial representations. In comparisons under identical memory constraints like AbsGS-0008, GDAGS exhibits full superiority in performance. Even when competing against AbsGS-004 under larger memory budgets, GDAGS maintains comparable results. GDAGS employs direction-aware adaptive dynamic weighting for individual Gaussians, promoting effective densification while optimizing spatial representations.
Qualitative Results. We present novel view synthesis results of GDAGS and baseline methods in Figure 3, demonstrating that our approach achieves high-quality visual fidelity across all scenes while effectively mitigating localized blur and detail loss, such as vegetation under benches, uneven concrete pavements, and textural details on train surfaces. Comprehensive quantitative and qualitative results are presented in the Appendix C.
@@
4.3 Analysis
4.3.1 Ablation analysis
We conduct ablation experiments on the proposed weighting method and trained two additional models: GDAGS-S and GDAGS-C, which represent split gradient weighting and clone gradient weighting, respectively, for experimental comparison with the original 3DGS and complete GDAGS. The experimental results are shown in Table 2. Split gradient weighting can effectively optimize SSIM and LPIPS indicators and reduce memory costs. Clone gradient weighting can improve PSNR indicators but increase memory costs. Integrating split gradient weighting and clone gradient weighting can achieve a balance between performance and efficiency.
@@ 4.3.2 Parameter sensitivity analysis
We conduct a parameter sensitivity analysis on the hyperparameter $p$ to reflect the impact of parameters on the model. We set the parameters to 10, 15, 20, 25, and 30, respectively, and the specific data is shown in Figure 4. In the previous analysis, we explained that the hyperparameter $p$ controls the number of Gaussians selected for densification in adaptive density control, thereby indirectly controlling the total number of Gaussians. The increase of hyperparameter $p$ decreases the number of Gaussians used for densification, thereby reducing memory and lowering rendering performance. The reduction of hyperparameter $p$ increases the number of Gaussians used for densification, thereby increasing memory and improving rendering performance.
@@ 4.3.3 Densification process analysis
We quantified the Gaussian split/clone counts per densification round and the total Gaussian population dynamics during training (see Figure 5). The experimental results demonstrate distinct behavioral patterns: In cloning operations, all three methods exhibit consistent trends, with 3DGS achieving the highest clone count due to its standard threshold. AbsGS, employing twice the cloning threshold of 3DGS, yields the lowest clone numbers, whereas GDAGS suppresses directionally inconsistent clones through gradient alignment constraints, resulting in intermediate cloning activity. GDAGS further ensures localized replenishment of directionally consistent clones while mitigating redundant accumulation of geometrically conflicting clones. For splitting operations, AbsGS generates excessive splits by indiscriminately utilizing gradient magnitudes, whereas 3DGS underperforms due to gradient direction conflicts. GDAGS effectively filters splittable Gaussians through adaptive criteria. Regarding total Gaussian counts, GDAGS exhibits a more stable trajectory and faster convergence compared to 3DGS and AbsGS, validating its optimized densification strategy in balancing geometric fidelity and computational efficiency.
@@ 5 Conclusion
This paper proposes GDAGS, a dynamic adaptive Gaussian splatting framework that addresses critical challenges in 3D scene reconstruction, including over-densification, over-reconstruction, and memory inefficiency. By introducing the directional consistency metric (GCR) and nonlinear dynamic weighting, GDAGS enables direction-aware densification control: GCR discriminates between directionally consistent/divergent Gaussians, while dynamic weights independently regulate densification at the per-Gaussian level. This dual mechanism effectively suppresses redundant Gaussian proliferation, eliminates outlier distributions, and drives compact spatial representations without relying on heuristic threshold tuning.
Appendix A Algorithm and weighted response curve
@@
3:Initialize gradient buffers $\{\nabla_{i,v}\}\leftarrow 0$
@@
11:   Compute $\mathcal{C}_{i}$ via Eq. (5)
12:   Calculate $w_{i}$ using Eq. (12)
@@
Appendix B Gaussian densification
Figure 6: Corresponding curves of weights for different hyperparameters pp.
Figure 7: Comparison of more detailed methods in the Gaussian densification process.
@@ Appendix B Gaussian densification
Figure 8: Visualization of different densification methods during the training process in flowers sense.
Appendix C Detailed results

### v2

@@ 4.1 Setup
Plenoxels 0.626 23.08 0.463 2.1GB 0.719 21.08 0.379 2.3GB 0.795 23.06 0.510 2.7GB
INGP 0.671 25.30 0.371 13MB 0.723 21.72 0.330 13MB 0.797 23.62 0.423 13MB
Mip-NeRF360 0.792 27.69 0.237 8.6MB 0.759 22.22 0.257 8.6MB 0.901 29.40 0.245 8.6MB
[…]
@@ 4.2 Comparison
Quantitative Results. We report the quantitative results of the experiment in the Table 1, which indicate that our method produces comparable results in most cases. It is worth noting that our method has less memory at performance close to or better than SOTA. Compared to Pixel-GS, GDAGS achieves comparable or superior rendering performance with only 20%-50% of the memory consumption. This, combined with its comparison against 3DGS, demonstrates that GDAGS effectively addresses over-densification through nonlinear dynamic weights that suppress redundant Gaussians during densification. By adaptively controlling density, GDAGS enables scene Gaussians to converge toward compact spatial representations.
Compared to AbsGS, which enforces non-negative gradient components to resolve directional conflicts, this approach introduces a new issue: it amplifies the influence of outlier gradients arising from noisy or suboptimal regions. As a result, it triggers excessive and often unnecessary splitting, leading to over-densification, higher memory consumption, and noisier renderings—without a proportional improvement in visual quality. In contrast, Pixel-GS employs a pixel-coverage weighting mechanism that enhances spatial adaptation. However, it fails to account for directional information in gradients. This omission means it still encourages densification in areas with large yet directionally consistent gradients, resulting in redundant Gaussians and elevated memory costs. Our method addresses these limitations by leveraging directional coherence to actively suppress such redundant densification.
Qualitative Results. We present NVS results of GDAGS and baseline methods in Figure 3, demonstrating that our approach achieves high-quality visual fidelity across all scenes while effectively mitigating localized blur and detail loss. Comprehensive quantitative and qualitative results are presented in the Appendix E.
4.3 Analysis
@@
[…]
4.3.1 Ablation study
We perform ablation studies on the proposed weighting strategy by training three model variants: GDAGS-S (applying weighting only during splitting), GDAGS-C (only during cloning), and GDAGS-L (using a linear weighting function $w_{i}=2-\mathcal{C}_{i}$). These are compared against the original 3DGS and the full GDAGS model. Results are summarized in Table 2. The study reveals that split-oriented weighting (GDAGS-S) effectively improves SSIM and LPIPS while reducing memory usage. Clone-oriented weighting (GDAGS-C) enhances reconstruction fidelity (PSNR) but at the cost of increased memory consumption. Integrating both strategies in the full GDAGS model achieves an optimal balance between performance and efficiency. Notably, the nonlinear weighting function in GDAGS significantly outperforms its linear counterpart (GDAGS-L), as it more effectively translates gradient direction coherence into adaptive control signals. This nonlinear mapping provides superior sensitivity to directional conflicts and consistency, leading to more stable optimization and higher-quality results across all metrics compared to the linear approximation.
@@ 4.3.2 Parameter sensitivity analysis
A sensitivity analysis is performed on hyperparameters $p$ and $\beta$, with detailed results for $\beta$ included in the Appendix D. The study shows that $p$ exerts a stronger influence on both rendering performance and computational efficiency. As demonstrated in Figure 4, we test $p$ across a range of values: 10, 15, 20, 25, and 30. As previously explained, $p$ determines the proportion of Gaussians selected for densification within the adaptive control mechanism. Increasing $p$ restricts densification to fewer Gaussians, which reduces memory usage but also compromises rendering quality. In contrast, a smaller $p$ expands the candidate set for densification, enhancing visual performance at the expense of greater memory overhead. These findings underscore the critical role of $p$ in balancing the trade-off between efficiency and reconstruction quality.
@@ 4.3.3 Densification process analysis
We quantified the Gaussian split/clone counts per densification round and the total Gaussian population dynamics during training in Figure 5. The experimental results demonstrate distinct behavioral patterns: In cloning operations, all three methods exhibit consistent trends, with 3DGS achieving the highest clone count due to its standard threshold. AbsGS, employing twice the cloning threshold of 3DGS, yields the lowest clone numbers, whereas GDAGS suppresses directionally inconsistent clones through gradient alignment constraints, resulting in intermediate cloning activity. GDAGS further ensures localized replenishment of directionally consistent clones while mitigating redundant accumulation of geometrically conflicting clones. For splitting operations, AbsGS generates excessive splits by indiscriminately utilizing gradient magnitudes, whereas 3DGS underperforms due to gradient direction conflicts. GDAGS effectively filters splittable Gaussians through adaptive criteria. Regarding total Gaussian counts, GDAGS exhibits a more stable trajectory and faster convergence compared to 3DGS and AbsGS, validating its optimized densification strategy in balancing geometric fidelity and computational efficiency.
4.3.4 Generalizability analysis
To demonstrate the generalizability of the proposed densification control strategy, we generalize our GDAGS module to two other advanced 3DGS frameworks: MCMC-3DGS (Kheradmand et al., 2024) and Compact-3DGS (Lee et al., 2024). As quantitatively shown in Table 3, the integration of GDAGS consistently improves the performance of both baselines, yielding superior results in terms of SSIM and LPIPS. This demonstrates the broad applicability and effectiveness of our gradient-direction-aware densification control in enhancing reconstruction quality across different 3DGS architectures. As shown in the qualitative results of Figure 6, the model integrated with GDAGS performs better on object surfaces and geometric edges, such as room floors, wardrobe gaps, etc.
[…]
Datasets Mip-NeRF360 Tanks&Temples Deep Blending
[…]
Figure 6: Qualitative analysis of GDAGS integrated in MCMC-3DGS and Compact-3DGS.
@@ 5 Conclusion
This paper proposes GDAGS, a dynamic adaptive Gaussian splatting framework that addresses critical challenges in 3D scene reconstruction, including over-densification, over-reconstruction, and memory inefficiency. By introducing the directional consistency metric (GCR) and nonlinear dynamic weighting, GDAGS enables direction-aware densification control: GCR discriminates between directionally consistent/divergent Gaussians, while dynamic weights independently regulate densification at the per-Gaussian level. This dual mechanism effectively suppresses redundant Gaussian proliferation, eliminates outlier distributions, and drives compact spatial representations without introducing additional heuristic densification thresholds.
ETHICS STATEMENT
This work adheres to the ICLR Code of Ethics. The study does not involve any human subjects or animal experimentation. All datasets used are publicly available and were processed in compliance with their respective licenses and applicable regulations concerning data privacy. No personally identifiable information was used, and no aspect of the study raises ethical, privacy, or security concerns. We are committed to maintaining transparency and scientific integrity in all stages of this research.
Reproducibility Statement
To ensure reproducibility, we have open-sourced the complete codebase, pre-trained models, and detailed experimental configurations at: https://anonymous.4open.science/r/GDAGS-D473. All datasets used are publicly available in Section 4. We provide an environment configuration file (environment.yml) and scripts to replicate the main results. All hyperparameters, random seeds, and hardware specifications are thoroughly documented.
Appendix A LLM USAGE
Large language models (LLMs) were used to assist in the writing and polishing of this manuscript. Specifically, an LLM was employed to help refine language expression, improve readability, and enhance clarity in certain sections of the paper. Tasks included sentence rephrasing, grammar checking, and improving the overall flow of the text. It is important to note that the LLM was not involved in the conception of ideas, research methodology, experimental design, or data analysis. All scientific content, conceptual development, and analytical results were generated solely by the authors. The use of the LLM was strictly limited to improving the linguistic quality of the manuscript and did not extend to any scientific reasoning or interpretation. The authors take full responsibility for the entire content of the manuscript, including any text modified or suggested by the LLM. We have carefully reviewed all AI-assisted content to ensure it adheres to ethical guidelines and accurately reflects the original intent and rigor of our research.
Appendix B Algorithm
@@
3:Initialize gradient buffers $\{\nabla_{i,v}\}\leftarrow 0$, indicate the gradient of the $i$-th Gaussian at view angle $v$
@@
11:   Compute $\mathcal{C}_{i}$ via Eq. equation 5
12:   Calculate $w_{i}$ using Eq. equation 6
@@
Appendix C Weight Function Response Curve Analysis
As shown in Figure 7, we systematically analyzed the influence of parameters $p$ and $\beta$ in the weight function. When $\beta=25$ is fixed, increasing the value of $p$ will cause the weight curve to be in $\mathcal{C}_{i}$. When the value is smaller, it becomes steeper, which means stronger punishment is imposed on Gaussian beams with inconsistent directions. When $p=15$ is fixed, $\beta$ mainly controls the overall magnitude of the weight function. These observation results are consistent with the quantitative analysis in the main text, verifying the rationality of parameter design.
Figure 7: Corresponding curves of weights for different hyperparameters $p$ and $\beta$.
Appendix D analyze
D.1 Hyperparameter sensitivity analysis
The hyperparameters beta were set to 15, 20, 25, 30, and 35 for sensitivity analysis, and the results are shown in Figure 8. The results show that the setting of beta has a relatively small impact on overall performance, and increasing beta can slightly improve performance and slightly increase the required storage space.
Figure 8: Performance of different hyperparameters $\beta$ in multiple datasets.
D.2 Densification process analysis
Figure 9 quantifies the Gaussian split/clone counts per densification round and the total Gaussian population dynamics during training. The results reveal distinct behavioral patterns across methods: In cloning, all three methods show consistent trends. 3DGS produces the most clones under its standard threshold, while AbsGS yields the fewest due to its doubled threshold. GDAGS suppresses directionally inconsistent clones via gradient alignment, resulting in intermediate numbers and ensuring localized replenishment without redundant accumulation. For splitting, AbsGS over-splits by using gradient magnitudes indiscriminately, whereas 3DGS under-splits due to gradient conflicts. GDAGS effectively filters splittable Gaussians with adaptive criteria. Overall, GDAGS exhibits a more stable trajectory and faster convergence in total Gaussian counts than 3DGS and AbsGS, confirming its balanced densification strategy between geometric fidelity and efficiency.
@@ D.2 Densification process analysis
Figure 9: Visualization of different densification methods during the training process in the flowers sense.
Table 4: Comparison of training time between power function and exponential function. The best result is bold.
Methods Mip-NeRF360 Tanks&Temples Deep Blending
GDAGS 1140s 555s 898s
GDAGS-E 1200s (+5.2%) 612s (+10.3%) 976s (+8.7%)
Table 5: Comparison of performance between power function and exponential function. The best result is bold.
Datasets Mip-NeRF360 Tanks&Temples Deep Blending
[…]
GDAGS 0.839 28.02 0.145 0.852 23.79 0.165 0.905 29.70 0.235
GDAGS-E 0.837 27.96 0.146 0.851 23.62 0.165 0.903 29.60 0.237
D.3 Nonlinear Dynamic Weighted Functions Analysis
Nonlinear functions are usually represented as power functions ($(1-\mathcal{C}_{i})^{p}$) and exponential functions ($e^{-p\mathcal{C}_{i}}$). GDAGS chooses to use the power function form for two main considerations. Firstly, from the perspective of computational load, the calculation of power functions is very efficient, while the calculation of exponential functions is more complex. The running time of torch.exp in PyTorch environment is longer than torch.pow. We construct a model consisting of exponential functions and named it GDAGS-E. We conduct additional exponential function ablation experiments, the training time and performance are reported in the Table 4 and Table 5. The results show that using exponential functions increases training time by 5%-10% compared to power functions.
Then, we analyze the properties of exponential and power functions themselves. We consider two candidate forms for the nonlinear weighting component, power function $f_{p}(x)=(1-x)^{p}$ and exponential function $f_{e}(x)=e^{-px}$. To compare their sensitivity characteristics, we analyze the absolute values of their derivatives: $|f_{p}^{\prime}(x)|=p(1-x)^{p-1}$ and $|f_{e}^{\prime}(x)|=pe^{-px}$. These derivatives represent the rate at which each weighting function responds to changes in gradient coherence. For a power function, it is a monotonically decreasing function over the domain $x\in[0,1]$, with a value range of $f_{p}\in\left[0,1\right]$. For a exponential function, it is a monotonically decreasing function over the domain $x\in[0,1]$, with a value range of $f_{p}\in\left[pe^{-p},1\right]$. We let $f(x)=(1-x)^{p}-e^{-px}$, whose derivative is $f^{\prime}=p[e^{-px}-x^{p-1}]$, and the corresponding curve of the function is shown in the Figure 10. In the domain, $f(0)=0$, $f(1)>0$, the derivative is first less than 0 and then greater than 0, and the zero point $x=x_{0}$ keeps approaching 0. This indicates that the power function curve is always below the exponential function curve and as $p$ increases, the exponential function continues to approach the power function. When $0<x<x_{0}$, $f^{\prime}<0$ indicates that the power function decreases faster than the exponential function in the $0,x_{0}]$ interval. When $x_{0}<x<1$, $f^{\prime}>0$ indicates that the power function decreases more smoothly than the exponential function in the $x_{0},1]$ interval. This is consistent with our design goal of selecting only Gaussians with significant gradient direction conflicts and maintaining consistent suppression for Gaussians with consistent gradient directions.
Figure 10: The function curve of the difference between power function and exponential function.
D.4 Analysis of Abnormal Behavior of Different Models on Dataset Deep Blending
We observed that the performance of GDAGS on the Deep Blending dataset is slightly lower than that of Mini plating, and the number of Gaussians in Mini plating is significantly less than that of GDAGS. The performance metrics reconstructed on other datasets show the same trend as the Gaussian number, while the opposite trend is observed on the Deep Blending dataset. Based on this phenomenon, we propose a hypothesis that the relatively simple geometric shapes and textures of the Deep Blending dataset make them more prone to overfitting when modeling with excessive Gaussian numbers. To test this hypothesis, we introduced a Gaussian dropout mechanism into GDAGS during training, creating variants dubbed GDAGS-DROP (random dropout) and GDAGS-ODROP (opacity-weighted dropout). The results are reported in Table 6. The performance of GDAGS-DROP initially improves with a moderate dropout rate before declining, which confirms that overfitting is a relevant factor on this dataset. The optimally regularized variant, GDAGS-ODROP-5%, surpasses all Mini-splatting configurations in SSIM and PSNR. At the same time, we consulted some 3DGS related work, and one of the papers, ControlGS (Zhang et al., 2025), studied this situation. ControlGS proposes that as the number of Gaussians increases, scene reconstruction will go through four stages: underfitting, effective state, saturation, and overfitting. ControlGS also uses pruning to keep the model in an effective state. Notably, ControlGS exhibits a similar cross-dataset performance pattern: it underperforms against other baselines on complex datasets like Mip-NeRF-360 and Tanks&Temples but excels on Deep Blending, mirroring our findings.
Table 6: Performance comparison of GDAGS-DROP with Mini-splating-P (partial pruning) and Mini-splatting-D (only perform densification without pruning) on the Deep Blending dataset.
Methods Gaussians (M) SSIM PSNR LPIPS
Mini-splatting 0.35 0.908 29.98 0.253
Mini-splatting-P 1.55 0.906 29.92 0.230
Mini-splatting-D 4.63 0.906 29.88 0.211
GDAGS (GDAGS-DROP-0%) 1.71 0.905 29.70 0.235
GDAGS-ODROP-5% 0.74 0.909 30.03 0.241
GDAGS-DROP-5% 0.74 0.908 29.91 0.242
GDAGS-DROP-10% 0.58 0.908 29.88 0.247
GDAGS-DROP-20% 0.32 0.906 29.78 0.255
D.5 Gradient Sparse Scene Analysis
To analyze the behavior of GDAGS and baseline models in sparse gradient regions, we trained GDAGS-H, 3DGS-H, AbsGS-H, and Pixel-GS-H models respectively, where - H represents using 50% of the training views to increase the range of sparse gradient regions. As shown in Figure 11, GDAGS has fewer artifacts and better geometric edges in gradient sparse areas such as the sky, walls, and ground. Although the assumption of GDAGS is that Gaussian needs to receive sufficient gradient information, GDAGS still works well in sparse gradient regions.
Figure 11: Qualitative results of GDAGS and baseline models in sparse views.
D.6 Analysis and Guidance of Hyperparameters
First, we emphasize that all experimental results are obtained using a single, fixed set of hyperparameters ($\alpha=0.8$, $\beta=25$, $p=15$) across all scenes and datasets, without any scene-specific tuning. The consistent improvements in rendering quality (PSNR, SSIM, LPIPS) and memory efficiency demonstrate the strong generalization capability and robustness of GDAGS under a fixed configuration.
Second, the roles of $\alpha$, $\beta$, and $p$ are well-defined and interpretable:
  1. 1.
    When the GCR approaches 1, the weight converges to $\alpha$. Since Gaussians with high GCR typically exhibit large gradient magnitudes, $\alpha$ (set between 0 and 1) acts as a suppression factor to prevent over-densification in well-reconstructed regions.
  2. 2.
    When GCR approaches 0, the weight becomes $\alpha+\beta$. These Gaussians often have small gradient norms due to directional conflicts. To promote their densification, we set $\beta$ to match the max_screen_size value from the original 3DGS implementation, ensuring compatibility and intuitive scaling.
  3. 3.
    The exponent $p$ controls the number of Gaussians selected for densification. Its effect is predictable: increasing $p$ restricts densification to fewer Gaussians, reducing memory usage at a potential cost to reconstruction quality; decreasing $p$ has the opposite effect. In practice, $\alpha$ and $\beta$ can remain unchanged with minimal impact on performance, and the value of $p$ can be adjusted according to different requirements (high rendering accuracy or low memory consumption).
Furthermore, our sensitivity analysis confirms that GDAGS maintains stable performance across a wide range of $p$ (10–30) and $\beta$ (15–35) values, reinforcing its generalizability without demanding extensive hyperparameter search.
Table 7: Efficiency analysis of GDAGS and baseline models. TT(s) in the table represents training time (in seconds). The best score is red.
Datasets Mip-NeRF360 Tanks&Temples Deep Blending
[…]
3DGS 1594 134 912 154 1344 137
Pixel-GS 1811 89 1450 92 1640 90
AbsGS 1571 111 621 125 937 142
[…]
D.7 Efficiency Analysis
We report the training time and inference time (FPS) of GDAGS compared to other baseline models in Table 7. As the results demonstrate, GDAGS achieves the fastest training convergence across all datasets. This efficiency stems from our gradient-direction-aware densification control, which strategically prioritizes the most impactful splitting and cloning operations. By avoiding redundant densification steps that contribute little to reconstruction quality, GDAGS not only produces a more compact scene representation but also reduces the total computational cost of training. Regarding the computational overhead of GCR, we argue it is negligible for two key reasons. First, the calculation of the GCR is computationally lightweight, involving only simple vector norms and summations on pre-existing gradients. Second, and more importantly, the GCR-driven densification control leads to a far more compact scene representation, often reducing the final number of Gaussians by hundreds of thousands to millions compared to 3DGS. This reduction directly accelerates subsequent optimization steps.
Appendix E Detailed results